OpenStudy (anonymous):

Standard normal distribution z-score is out of range, what do I do? I have a statistics problem that we have to find the probability of something. The mean=64.3, standard deviation=2.6, sample size=60 and what is the probability if the mean height of the sample is greater than 66?

OpenStudy (anonymous):

So I did this: \[Probability = 1 - z=(\frac{ 66-64.3 }{ 2.6/\sqrt{60} })=1- z(\frac{ 1.7 }{ 0.33566 })=1-z(5.06465)\] What do I do to find the z-score? I know probability can be found for anything of z-score with a range of \[-\infty,\infty\] But my z-score standard normal distribution chart only goes up from -3.4 to 3.4. Thanks!

OpenStudy (freckles):

yeah shouldn't F(5.06) be close to 1 that is what the values on your chart approach

OpenStudy (freckles):

so if you have everything right above probability is approximate 1-1